# DeltaPhi: Learning Physical Trajectory Residual for PDE Solving

### v1

@@ B.1.1 Regular Domain
Darcy Flow. Darcy Flow is a steady-state solving problem. We conduct the experiment on Darcy Flow using the same dataset as [2], consisting of $421\times 421$ resolution fields with Dirichlet boundary. The low resolution data e.g. $22\times 2$ are obtained via uniformly downsampling operation.
Navier-Stokes. Navier-Stokes is a challenging time-series solving problem. We use the public dataset from  [2]. The viscosity of trajectories is $51e-$ 5 ($Re=2000$). The spatial resolution is $64\times 64$. The input and output time step length are both $10$.
@@ B.1.2 Irregular Domain
For these irregular domain problems, we take the exactly same setting with NORM, more details about the dataset can be found in [10].
@@ B.2 Base Model
Clifford Fourier Neural Operator (CFNO) [16]. Clifford Fourier Neural Operator (CFNO) [16] employs the Clifford Algebra in the neural network architecture, incorporating geometry prior between multiple physical fields. We employ the official implementation for the model architecture. The signature is set as $(-1,-1)$ and we pad lacked physical channels with zero. All hyperparameters except the channel width (32 in CFNO) remain consistent with FNO. The training setup (including the optimizer and scheduler) is the same as FNO.
@@ B.3 Implementation Details
Except for specific statements, the experimental details are as follows:
  • Retrieval: For training, the auxiliary trajectory retrieval range $K$ is set as 20. During inference, the auxiliary trajectory with the most similarity with $a_{i}$ is retrieved.
  • Input: The customized input function is partial (last 3 channels) $a_{k_{i}}$ and complete $u_{k_{i}}$.
  • Optimizing: We use the Adam [46] optimizer with initial learning rate $0.001$ and weight decay $0.0001$. The StepLR scheduler with step size 100 and $\gamma=0.5$ is used to control the learning rate. The batch size is set as 8. All models are trained for 500 epochs. All experiments could be conducted on a single NVIDIA GeForce RTX 4090 device.
@@ B.4 Metric
where $\hat{u}_{i}$ and $u_{i}$ are predicted trajectory and real trajectory respectively.
@@ C.1 Selection of $K$ Value
For more challenging generalization problems (eg. serious data bias problems in realistic scenarios), we provide an empirical selection algorithm for the "initial value" decision of $K$ (the "final value" of $K$ is also related to the base model’s capability). Given the training set (noted as $\{a^{train}_{i}\}_{i=1}^{N_{train}}$), suppose the input functions (noted as ${\{a^{test}_{i}\}}_{i=1}^{N_{test}}$) of pending solved physical trajectories is available. The initial value of $K$ could be calculated with the following steps: (1) Retrieve the most similar trajectory (noted as $\{a^{test}_{k_{i}}\}_{i=1}^{N_{test}}$) from the training set for every pending solving function. We denote the similarity values between $a^{test}_{i}$ and $a^{test}_{k_{i}}$ as $s^{test}_{i}$. (2) Calculate the maximum value $s^{test}_{max}$ of $\{s^{test}_{i}\}_{i=1}^{N}$. (3) For every training trajectory $a^{train}_{i}$, calculate its similarity with other training trajectories. Denote the index of $r$-th similar trajectory with $a^{train}_{i}$ as $k^{r}_{i}$, and the similar score between $a^{train}_{i}$ and $a^{train}_{k^{r}_{i}}$ as $s^{r}_{i}$. (4) The $K$ is calculated as the minimum $r$ satisfying $s^{test}_{max}\leq s_{i}^{r}$ for any $i\in[1,N^{train}]$.
C.2 Variant Resolution Auxiliary Trajectory Integration
Sometimes the auxiliary function $a_{k_{i}}$ has a different resolution. For example, when we need to predict high-resolution physical trajectory using low-resolution trained models, the auxiliary samples $a_{k_{i}}$ in the training set are more coarse-grained than the queried function $a^{H}$. In this scenario, we could interpolate the low-resolution auxiliary sample to a fine-grained grid, then concatenate the interpolated high-resolution grid with the queried function $a^{H}_{i}$.
$a_{input}=\mathcal{F}^{inter}(a_{k_{i}})\oplus a^{H}.$ (13)
In this work, we implement the interpolation method $\mathcal{F}^{inter}$ based on the Fourier Transform. Specifically, we first transform the coarse-grained input $u_{k_{i}}$ into Fourier Space, fill zeros in the high-frequency region to target resolution, and finally transform it back to spatial space. Our experimental results show the trained residual neural operator could generalize well in this interpolation scenario.
C.3 Variant Resolution Trajectory Retrieval
For the high resolution inference scenario, we could first downsample the high-resolution queried function $a^{H}$ to low resolution $a^{H}_{low}$, then calculate the similarity score:
$\mathcal{F}^{ret}(a^{H},\mathcal{T})$ $=\text{argmax}_{j}\,\text{Sim}(a^{H}_{low},a_{j}),$ (14)
$a^{H}_{low}$ $=\mathcal{F}^{low}(a^{H}),$
where $\mathcal{F}^{low}$ is down sampling function. In this work, we implement $\mathcal{F}^{low}$ by erasing additional information in high-frequency regions based on the Fourier Transform.
C.4 Complexity Analysis of Physical Trajectory Retrieval
@@
The computational cost for retrieval is negligible compared to neural network inference. Here is a detailed analysis of the computational cost for physical trajectory retrieval.
@@
The computational expensive is mainly related to two factors: the training set size (noted as $N_{train}$) and the number of physical field points (ie. the resolution of physical field) (noted as $N_{field}$). The computing process consists of the following two steps:
@@
  • Step 1: Similarity Score Calculation. For every sample in the retrieval set, the similarity metric function (eg. Cosine Similarity) could be calculated with linear time complexity $O(N_{field})$. Thus the time complexity of calculating similarity scores for all retrieved samples is $O(N_{field}\cdot N_{train})$.
@@
  • Step 2: Ranking based on similarity scores. We use the simple sorting algorithm (Quicksort) to rank all samples based on their similarity score. The time complexity is $O(N_{train}\cdot\log N_{train})$. The overall time complexity for retrieval is $O(N_{field}\cdot N_{train}+N_{train}\cdot\log N_{train})$.
Commonly, the numerical ranges for $N_{field}$ and $N_{train}$ are approximately between $10^{3}$ to $10^{4}$, and $10^{2}$ to $10^{4}$, respectively. Therefore, the real computational amount is about $O(10^{8})$, this could be calculated on most computers. In our implementation, the average retrieval time on Darcy Flow ($85\times 85$ resolution) is 2.67e-3 to 1.84e-2 seconds for different retrieval set sizes (100 to 900).
@@ D.1 Similarity between $u(x)$ Visulization Detail
We visualize the correlation between $u(x)$ normalized distance and $a(x)$ similarity rank (shown in Figure 4) as following steps: Firstly, taking the training set $\mathcal{T}$ as retrieval set, we sample auxiliary trajectory $(a_{k_{i,r}},u_{k_{i,r}})$ with $r$-th top similarity score for every test sample $(a^{test}_{i},u^{test}_{i})$ in the test set. Here $r$ represents the similarity rank (higher $r$ indicates lower similarity), and we take gradually increasing values from $1$ to $80$ for $r$. Next, we calculate the normalized distance between $u^{test}_{i}$ with its auxiliary sample $u_{k_{i,r}}$. The normalized distance is defined the same as the relative error Equation 14.
@@ D.2 Label Distribution Visualization Detail
We visualize the label distribution (as shown in  Figure 5) by reducing the high-dimension function fields to 2 dimensions utilizing Principle Component Analysis. Specifically, for both direct operator mapping and residual operator mapping, we first optimize a PCA model on the training labels (100 labels) and then calculate the dimension-reduced labels for every trajectory in the training and testing set. Finally, we visualize every two-dimension label point on the graph as well as the ellipses representing the stand deviation and the range along each dimension.

### v2

@@ B.1.1 Regular Domain
Darcy Flow. Darcy Flow is a steady-state solving problem. We conduct the experiment on Darcy Flow using the same dataset as [2], consisting of $421\times 421$ resolution fields with Dirichlet boundary. The low resolution data e.g. $22\times 22$ are obtained via uniformly downsampling operation.
Navier-Stokes. Navier-Stokes is a challenging time-series solving problem. We use the public dataset from  [2]. The viscosity of trajectories is 1e51e-5 ($Re=2000$). The spatial resolution is $64\times 64$. The input and output time step length are both $10$.
@@ B.1.2 Irregular Domain
For these irregular domain problems, we take the exact same setting with NORM, more details about the dataset can be found in [9].
@@ B.2 Base Model
Clifford Fourier Neural Operator (CFNO) [11]. Clifford Fourier Neural Operator (CFNO) [11] employs the Clifford Algebra in the neural network architecture, incorporating geometry prior between multiple physical fields. We employ the official implementation for the model architecture. The signature is set as (1,1)(-1,-1) and we pad lacked physical channels with zero. All hyperparameters except the channel width (32 in CFNO) remain consistent with FNO. The training setup (including the optimizer and scheduler) is the same as FNO.
@@ B.3 Implementation Details
Except for specific statements, the experimental details are as follows: (a) Retrieval: For training, the auxiliary sample retrieval range $K$ is set as 20. During inference, the auxiliary sample with the most similarity with $a_{i}$ is retrieved. (b) Input: The customized input function is partial (last 3 channels) $a_{k_{i}}$ and complete $u_{k_{i}}$. (c) Optimizing: We use the Adam [57] optimizer with initial learning rate $0.001$ and weight decay $0.0001$. The StepLR scheduler with step size 100 and $\gamma=0.5$ is used to control the learning rate. The batch size is set as 8. All models are trained for 500 epochs. All experiments could be conducted on a single NVIDIA GeForce RTX 4090 device.
@@ B.4 Metric
where $\hat{u}_{i}$ and $u_{i}$ are predicted solution and real solution respectively.
@@ C.1 Selection of $K$ Value
For more challenging generalization problems (eg. serious data bias problems in realistic scenarios), we provide an empirical selection algorithm for the "initial value" decision of $K$ (the "final value" of $K$ is also related to the base model’s capability). Given the training set (noted as $\{a^{train}_{i}\}_{i=1}^{N_{train}}$), suppose the input functions (noted as ${\{a^{test}_{i}\}}_{i=1}^{N_{test}}$) of pending solved physical trajectories is available. The initial value of $K$ could be calculated with the following steps: (1) Retrieve the most similar sample (noted as $\{a^{test}_{k_{i}}\}_{i=1}^{N_{test}}$) from the training set for every pending solving function. We denote the similarity values between $a^{test}_{i}$ and $a^{test}_{k_{i}}$ as $s^{test}_{i}$. (2) Calculate the maximum value $s^{test}_{max}$ of $\{s^{test}_{i}\}_{i=1}^{N}$. (3) For every training sample $a^{train}_{i}$, calculate its similarity with other training trajectories. Denote the index of $r$-th similar sample with $a^{train}_{i}$ as $k^{r}_{i}$, and the similar score between $a^{train}_{i}$ and $a^{train}_{k^{r}_{i}}$ as $s^{r}_{i}$. (4) The $K$ is calculated as the minimum $r$ satisfying $s^{test}_{max}\leq s_{i}^{r}$ for any $i\in[1,N^{train}]$.
C.2 Complexity Analysis of Auxiliary Sample Retrieval
@@
The computational cost for retrieval is negligible compared to neural network inference. Here we provide both theoretical complexity analysis and comprehensive experimental validation.
@@
Theoretical Analysis. The computational complexity primarily depends on two factors: the training set size (denoted as $N_{train}$) and the number of physical field points (i.e., the resolution of physical field) (denoted as $N_{field}$). The retrieval process consists of two main steps:
@@
  • Step 1: Similarity Score Calculation. For each sample in the retrieval set, computing the similarity metric (e.g., Cosine Similarity) has linear time complexity $O(N_{field})$. Thus, calculating similarity scores for all retrieved samples has complexity $O(N_{field}\cdot N_{train})$.
@@
  • Step 2: Similarity Score Ranking. Using QuickSort to rank samples based on similarity scores has complexity $O(N_{train}\cdot\log N_{train})$.
The overall time complexity is $O(N_{field}\cdot N_{train}+N_{train}\cdot\log N_{train})$. With typical values of $N_{field}$ and $N_{train}$ ranging from $10^{3}$ to $10^{4}$ and $10^{2}$ to $10^{4}$ respectively, the computation is feasible on standard hardware.
Experimental Validation. We conducted extensive experiments to measure the actual computational overhead of retrieval:
  • Retrieval Time Analysis: We tested retrieval time across different training set sizes on both CPU and GPU (single RTX 3090) for Darcy Flow ($85\times 85$ resolution):
    Table 15: Retrieval time (in seconds) for different training set sizes
    Device 100 300 500 700 900
    CPU 2.67e-3 5.57e-3 9.05e-3 1.32e-2 1.66e-2
    GPU 1.89e-4 1.83e-4 1.64e-4 2.43e-4 3.06e-4
  • End-to-End Inference Time: We compared the total inference time per example between direct learning and residual learning across different models:
    Table 16: Average inference time (in seconds) comparison
    Method FNO FFNO Galerkin
    Direct Learning 1.20e-3 2.39e-3 5.01e-3
    Residual Learning 1.67e-3 2.85e-3 5.58e-3
    Increased Time 4.67e-4 4.66e-4 5.75e-4
The experimental results demonstrate that:
  • GPU-based retrieval is highly efficient, taking only 0.2-0.3ms across different training set sizes.
  • The total additional time for residual learning, including retrieval and data preparation, is consistently around 0.5ms.
These comprehensive experiments confirm that the auxiliary sample retrieval introduces negligible computational overhead and does not significantly impact the practical deployment of our method.
C.3 Discussion on Boundary Conditions
In this work, DeltaPhi does not explicitly process boundary conditions as a separate input. Instead, it implicitly handles them by leveraging the full auxiliary solution $u_{k_{i}}$ as a strong physical prior. This auxiliary solution, which is a key input to the model, already satisfies the boundary conditions of its corresponding system. The model’s effectiveness in handling complex boundary effects is demonstrated by its strong performance on irregular domain problems, such as Pipe Turbulence and Blood Flow, which feature complex inlet/outlet conditions. While our current approach is effective, future work could explore enhancing the retrieval mechanism with a boundary-aware similarity metric for enhanced residual neural operator.
@@ D.1 Similarity between $u(x)$ Visulization Detail
We visualize the correlation between $u(x)$ normalized distance and $a(x)$ similarity rank (shown in Figure˜4) as following steps: Firstly, taking the training set $\mathcal{T}$ as retrieval set, we retrieve auxiliary sample $(a_{k_{i,r}},u_{k_{i,r}})$ with $r$-th top similarity score for every test sample $(a^{test}_{i},u^{test}_{i})$ in the test set. Here $r$ represents the similarity rank (higher $r$ indicates lower similarity), and we take gradually increasing values from $1$ to $80$ for $r$. Next, we calculate the normalized distance between $u^{test}_{i}$ with its auxiliary sample $u_{k_{i,r}}$. The normalized distance is defined the same as the relative error in Equation˜19.
@@ D.2 Label Distribution Visualization Detail
We visualize the label distribution (as shown in  Figure˜5) by reducing the high-dimension function fields to 2 dimensions utilizing Principle Component Analysis. Specifically, for both direct operator mapping and residual operator mapping, we first optimize a PCA model on the training labels (100 labels) and then calculate the dimension-reduced labels for every sample in the training and testing set. Finally, we visualize every two-dimension label point on the graph as well as the ellipses representing the stand deviation and the range along each dimension.